OpenStudy (anonymous):

A car is traveling at a constant velocity. After two hours the car has traveled 110 miles. After four hours the car has traveled 220 miles. Assuming a linear equation fits this situation, find the distance that the car will travel after five hours.

OpenStudy (anonymous):

110/2=55 2*55=110 220/4=55 4*55=220 5*55=........